OpenStudy (anonymous):

Is it possible that the car could have accelerated to 55mph within 268 meters if the car can only accelerate from 0 to 60 mph in 15 seconds?

OpenStudy (anonymous):

60 - 0/ 15 = 4 m/s^2

OpenStudy (lgbasallote):

depends what the initial velocity was...

OpenStudy (lgbasallote):

if initial velocity is 54 mph then it is very possible

OpenStudy (anonymous):

and the car starts from rest @lgbasallote

OpenStudy (lgbasallote):

where did it say that?

OpenStudy (lgbasallote):

as far as my English is concerned, i do not see "car starts from rest"

OpenStudy (anonymous):

probably...not possible..

OpenStudy (anonymous):

@lgbasallote u see 0 to 60 mph in 15 seconds?

OpenStudy (anonymous):

hm....i have done a mistake

OpenStudy (lgbasallote):

it CAN accelerate from 0 to 60 mph in 15 seconds

OpenStudy (anonymous):

i have nt changed the same

OpenStudy (anonymous):

hm.....

OpenStudy (anonymous):

i will rewrite

OpenStudy (lgbasallote):

but it doesnt say the ACTUAL intial velocity of the car

OpenStudy (anonymous):

First convert meter into miles

OpenStudy (anonymous):

The car can accelerate from 0 to in 15 seconds. Then maximum acceleration is 26.8/15 = 1.8m/s^2

OpenStudy (lgbasallote):

it was asking if it was possible if a car accelerated TO 55 mph that means there should be a from...so what is that from

OpenStudy (anonymous):

and...From ur question ..... V^2 - u^2 = 2as 24.6 * 24.6 / 2 *268 = 1.13m/s^2 = a

OpenStudy (anonymous):

This acceleration is less than the maximum possible acceleration, so the car can reach the speed 55 mph within 268 meters.

OpenStudy (anonymous):

@KarlaKalurky Did u understand...

OpenStudy (anonymous):

why did you divide 26.8 to 15s?

OpenStudy (anonymous):

1 miles = 1609.34meter

OpenStudy (anonymous):

why?